2-5. With u = 3.986 X 1014 mº/s?, a = 7.0 x 100 m, and € = 0.35: (a) Find E, H, Ip, la, Vp, and Va. (b) Find r, V, and d whenwhere a is half of the major axis of an elliptical orbit. This is an orbital mechanics problem
Answer #1

Answer:- Given that l= 31986* 1010 m3/s² a = 7,0x106m € = 0.35 (a) va = a (He) = 7:08 106 (1+0.35) = 7x106 (185) [vaa 9, 65 xva = (1-e) Va 29.2 = (1-0,35) 3,986x1014 9.651106 Vp Up, 38103 x 9.650106 4.558 106 Up = 8,05x103) (C) y = 8,813 x 106m coso0:35 Sin (150) 10:35 COS(150) 10135 Sin (150) 1+0-3500S(150) & tant (0.35 Tp = 14,090 6 31986 x 1014 8,8132106 2 ( 27.83X106